# Improving Cross-lingual Representation for Semantic Retrieval with Code-switching

### v1

@@ 1 Introduction
Specifically, mBERT (Devlin et al., 2019), XLM (Conneau and Lample, 2019), Unicoder (Huang et al., 2019), CMLM (Yang et al., 2021) and ERNIE-M (Ouyang et al., 2021) learn the cross-lingual sentence representation mainly using masked language modeling (MLM). For other objective functions, MMTE (Siddhant et al., 2020) exploits multi-lingual machine translation, and CRISS (Tran et al., 2020) leverages unsupervised parallel data mining. Some models use Siamese network architectures to better adapt them to SR. For example, InferSent (Conneau et al., 2017) uses natural language inference (NLI) datasets to train the Siamese network. USE (Yang et al., 2019), M-USE (Yang et al., 2020b) and LaBSE (Feng et al., 2022a) exploit ranking loss. SimCSE (Gao et al., 2021), InfoXLM (Chi et al., 2021) and HICTL (Wei et al., 2021) use contrastive learning. However, the previous highly similar approaches almost ignore the transmission of some features of the downstream tasks to PTMs. In other words, most methods (Chi et al., 2021; Luo et al., 2021) directly fine-tune the models on downstream tasks without providing any signals related to SR. In addition, they are mainly pre-trained on combined monolingual data where few of the sentences are code-switched. Since the user queries often contain many code-switched sentences, it is insufficient to exploit the commonly used methods directly for the SR task in the e-commerce scenario.
@@ 3.1 Alternative Cross-lingual PTM
The main architecture of our model is shown in 2. We jointly train the PTM on the code-switched data using the cross-lingual masked language model (XMLM) and the similarity loss to address the limitation of existing PTMs that are trained without signals directly related to the downstream tasks (e.g., SR). Contrarily, we add a similarity loss term to the pre-training objective to adjust the similarity between input ($query$) and output ($label$). Thus, the similarity between $query$ and $label$ has been controlled by exploiting the similarity loss during the continual pre-training step. For the sentence-level SR task, the given knowledge is composed of a certain number of $\langle query,label\rangle$ pairs (see Figure 2). We regard $\mathbf{q}$ as a $query$ and take $\mathbf{l}$ as a corresponding $label$. Given a query $\mathbf{q}=\{q_{1},\dots,q_{i},\dots,q_{I}\}$ and a label $\mathbf{l}=\{l_{1},\dots,l_{j},\dots,l_{J}\}$, the standard retrieval models usually formulate the sentence-level SR as a calculation of the similarity between $query$ and $label$ on the semantic space:
@@ 3.2 Building Code-switched Data for SR
Table 4: The comparison with accuracy score (Top-30 queries) between baseline systems on business corpora.
@@ 4.1 Setup
Figure 3: The comparison with average accuracy score (Top-10/20 queries) on the business dataset.
@@ 4.2 Main Results
Table 4 shows the retrieving results of the proposed method on Ali-Express, DARAZ, and LAZADA corpora by evaluating the accuracy on TOP-30 retrieved queries, respectively. The conduction of our experiment is composed of two steps, firstly we continually pre-train the models by combining the queries and labels among the training set. Then we conduct the fine-tuning on different languages by exploiting their own train set and dev set. Unlike other baselines, in our continual pre-training step, we utilize code-switched queries and labels instead of combining the original data to train our model. Among the baselines, VECO achieves better results on almost every language from the business corpus. The code-switching method has the most positive effects both on Id, Fil (The corpora of these two languages are highly code-switched. See Table 2) and Bn, Si (DARAZ), but brings fewer benefits for Zh (Ali-Express). As we do not leverage any code-switched data for En, we obtain less improvement than VECO. However, our approach consistently outperforms all the baselines on each language except En. As depicted in Figure 3, we also evaluate the models with accuracy scores on Top-10 and Top-20 queries (For the details, see Table 7 & 8 in Appendix).
@@
Table 5: The comparison with various evaluation metric on AskUbuntu corpus. “P@1" and “P@5" denote the precision score on Top-1 query and Top-5 queries, respectively. “Acc." represents the accuracy score on Top-30 queries.
@@
Table 6: The comparison with MRR@10 evaluation metric on the Quora Duplicate Questions dataset. The meaning of “Acc." is similar to that in Table 5.
Model MRR Acc.
@@
Figure 5: The effect of similarity loss $\mathcal{L}^{(\mathbf{q},\mathbf{l})}_{sim}$) and different values of the hyper-parameters $\lambda$ and Code-mixing Rate ($Cmd_{r}$) in our model on business corpora with accuracy score (Top-30 queries). (a) “w/" and “w/o" denote the accuracy score with or without similarity loss. (b) and (c) also represent the average accuracy score with different values of $\lambda$ (default value is $0.2$) and $Cmd_{r}$ (default value is $10\%$), respectively.
Verification of Robustness
As shown in Table 5, we also further explore the performance of our model on the openly available corpus AskUbuntu (Lei et al., 2016) and Tatoeba benchmark (Artetxe and Schwenk, 2019). In this experiment, we conduct continual pre-training and fine-tuning by leveraging only the AskUbuntu corpus without using other datasets. We evaluate all the baselines and our model using different evaluation metrics p@1, p@5, and accuracy with Top-1, Top-5, and Top-30 queries, respectively. Moreover, we further verify the effectiveness of our approach on another openly available benchmark Tatoeba and obtain remarkably better results than baselines. For more details, see Table 10 in Appendix.
4.3 Monolingual Semantic Retrieval
[…]
4.4 Ablation Study
The Effect of Similarity Loss $\mathcal{L}^{(\mathbf{q},\mathbf{l})}_{sim}$
The Effect of $\lambda$
The Effect of Code-switching
@@ Appendix A Case Study
Table 10: The comparison with accuracy score (Top-30 queries) on Tatoeba corpus for each language.
@@
Table 11: The effect of $\lambda$ on business corpora with accuracy score (Top-30 queries).
@@
Table 12: The effect of code-switching rate (“$Cmd_{r}$") on business corpora with accuracy score (Top-30 queries).
@@ Appendix B Results on Business Datasets
As shown both in Table 7 and Table 8, we make some explorations for the retrieving skill of our introduced approach on Ali-Express, DARAZ, and LAZADA corpora by evaluating the accuracy on TOP-10 and TOP-20 retrieved queries, respectively. The conduction of our experiment is composed of two steps: first, we continually pre-train the models by combining the queries and labels among the training set. Then we conduct the fine-tuning on different languages by exploiting their own train set &\&& dev set. Similar to the results on TOP-30 queries (see Table 4), VECO also obtains higher results among the baselines systems. However, our proposed model achieves remarkably better results than all baselines on each of the languages of the three business datasets.
@@ Appendix C Results on the Task of STS
[…]

### v2

@@ 1 Introduction
Specifically, mBERT (Devlin et al., 2019), XLM (Conneau and Lample, 2019), Unicoder (Huang et al., 2019), CMLM (Yang et al., 2021) and ERNIE-M (Ouyang et al., 2021) learn the cross-lingual sentence representation mainly using masked language modeling (MLM). For other objective functions, MMTE (Siddhant et al., 2020) exploits multi-lingual machine translation, and CRISS (Tran et al., 2020) leverages unsupervised parallel data mining. Some models use Siamese network architectures to adapt them to SR better. For example, InferSent (Conneau et al., 2017) uses natural language inference (NLI) datasets to train the Siamese network. USE (Yang et al., 2019), M-USE (Yang et al., 2020b) and LaBSE (Feng et al., 2022a) exploit ranking loss. SimCSE (Gao et al., 2021), InfoXLM (Chi et al., 2021) and HICTL (Wei et al., 2021) use contrastive learning. However, the previous highly similar approaches almost ignore the transmission of some features of the downstream tasks to PTMs. In other words, most methods (Chi et al., 2021; Luo et al., 2021) directly fine-tune the models on downstream tasks without providing any signals related to SR. In addition, they are mainly pre-trained on combined monolingual data where few of the sentences are code-switched. Since the user queries often contain many code-switched sentences, it is insufficient to exploit the commonly used methods directly for the SR task in the e-commerce scenario.
@@ 3.1 Alternative Cross-lingual PTM
The main architecture of our model is shown in Figure 2. We jointly train the PTM on the code-switched data using the cross-lingual masked language model (XMLM) and the similarity loss to address the limitation of existing PTMs that are trained without signals directly related to the downstream tasks (e.g., SR). Contrarily, we add a similarity loss term to the pre-training objective to adjust the similarity between input ($query$) and output ($label$). Thus, the similarity between $query$ and $label$ has been controlled by exploiting the similarity loss during the continual pre-training step. For the sentence-level SR task, the given knowledge is composed of a certain number of $\langle query,label\rangle$ pairs (see Figure 2). We regard $\mathbf{q}$ as a $query$ and take $\mathbf{l}$ as a corresponding $label$. Given a query $\mathbf{q}=\{q_{1},\dots,q_{i},\dots,q_{I}\}$ and a label $\mathbf{l}=\{l_{1},\dots,l_{j},\dots,l_{J}\}$, the standard retrieval models usually formulate the sentence-level SR as a calculation of the similarity between $query$ and $label$ on the semantic space:
@@ 3.2 Building Code-switched Data for SR
Table 4: The comparison with Recall@30 between baseline systems on business corpora.
@@ 4.2 Main Results
Table 4 shows the retrieving results of the proposed method on Ali-Express, DARAZ, and LAZADA corpora by evaluating the Recall@30 score, respectively. The conduction of our experiment is composed of two steps, firstly we continually pre-train the models by combining the queries and labels among the training set. Then we conduct the fine-tuning on different languages by exploiting their own train set and dev set. Unlike other baselines, in our continual pre-training step, we utilize code-switched queries and labels instead of combining the original data to train our model. Among the baselines, VECO achieves better results on almost every language from the business corpus. The code-switching method has the most positive effects both on Id, Fil (The corpora of these two languages are highly code-switched. See Table 2) and Bn, Si (DARAZ), but brings fewer benefits for Zh (Ali-Express). As we do not leverage any code-switched data for En, we obtain less improvement than VECO. However, our approach consistently outperforms all the baselines on each language except En. As depicted in Figure 3, we also evaluate the models with Recall@{10,20} (For the details, see Table 7 & 8 in Appendix).
Figure 3: The comparison with average Recall@{10,20} on the business dataset.
@@
Verification of Robustness
As shown in Table 5, we also further explore the performance of our model on the openly available corpus AskUbuntu (Lei et al., 2016) and Tatoeba benchmark (Artetxe and Schwenk, 2019). In this experiment, we conduct continual pre-training and fine-tuning by leveraging only the AskUbuntu corpus without using other datasets. We evaluate all the baselines and our model using different evaluation metrics P@1, P@5, and R@30 with Top-1, Top-5, and Top-30 queries, respectively. Moreover, we further verify the effectiveness of our approach on another openly available benchmark Tatoeba and obtain remarkably better results than baselines. For more details, see Table 10 in Appendix.
4.3 Monolingual Semantic Retrieval
As shown in Table 6, we also tend to verify the semantic retrieving skill of our approach on another openly available Quora Duplicate Questions dataset, which only includes English monolingual data. Since this corpus only provides the test set, we merge the English part of the train set from BUCC18 as our monolingual data. Then we continually pre-train each model to evaluate them on the Quora Duplicate Questions dataset without using the fine-tuned model. In this dataset, InfoXLMLarge obtains better results than other baselines, but our method outperforms all the baselines, which indicates that our approach has better retrieving skills compared to similar methods.
Table 5: The comparison with various evaluation metric on AskUbuntu corpus. “P@1" and “P@5" denote the precision score on Top-1 query and Top-5 queries, respectively. “R@30” denotes the recall score on Top-30 queries.
@@ 4.3 Monolingual Semantic Retrieval
4.4 Ablation Study
The Effect of Similarity Loss $\mathcal{L}^{(\mathbf{q},\mathbf{l})}_{sim}$
The Effect of $\lambda$
The Effect of Code-switching
Table 6: The comparison with MRR@10 and Recall@30 evaluation metrics on the Quora Duplicate Questions dataset.
Model MRR@10 Recall@30
@@ 4.4 Ablation Study
Figure 5: The effect of similarity loss $\mathcal{L}^{(\mathbf{q},\mathbf{l})}_{sim}$) and different values of the hyper-parameters $\lambda$ and Code-mixing Rate ($Cmd_{r}$) in our model on business corpora with Recall@30. (a) “w/" and “w/o" denote the accuracy score with or without similarity loss. (b) and (c) also represent the average Recall@30 score with different values of $\lambda$ (default value is $0.2$) and $Cmd_{r}$ (default value is $10\%$), respectively.
@@ 6 Conclusion and Future work
Acknowledgments
We are deeply thankful to each of our esteemed co-authors for their precious contributions and perceptive comments, which greatly improved the standard of this paper before its submission. Their knowledge and cooperative attitude were vital in forming our research. Moreover, our appreciation goes to our peers who carefully examined our work and identified minor mistakes, guaranteeing the precision and excellence of our conclusions. Their meticulous attention to detail substantially bolstered the credibility of our investigation.
@@ Appendix A Case Study
Table 10: The comparison with Recall@30 on Tatoeba corpus for each language.
@@
Table 11: The effect of $\lambda$ on business corpora with Recall@30.
@@
Table 12: The effect of code-switching rate (“$Cmd_{r}$") on business corpora with Recall@30.
@@ Appendix B Results on Business Datasets
As shown both in Table 7 and Table 8, we make some explorations for the retrieving skill of our introduced approach on Ali-Express, DARAZ, and LAZADA corpora by evaluating Recall@{10,20}, respectively. The conduction of our experiment is composed of two steps: first, we continually pre-train the models by combining the queries and labels among the training set. Then we conduct the fine-tuning on different languages by exploiting their own train set &\&& dev set. Similar to the results on TOP-30 queries (see Table 4), VECO also obtains higher results among the baselines systems. However, our proposed model achieves remarkably better results than all baselines on each of the languages of the three business datasets.
@@ Appendix C Results on the Task of STS
We regard that there exists a bit of difference between the two tasks, such as semantic retrieval (SR) and semantic textual similarity (STS). But both of them take sentence-level representation as a backbone. Therefore, we make some investigations on the task of STS. As shown in Table 9, we also make further validation on the similar task STS by comparing the semantic representation skill of the baseline models and our proposed method. In this experiment, all of the test sets (STS12, STS13, STS14, STS15, STS16, STS-B, and SICK-R) only include English sentences. Thus we evaluate the baselines and our model only using the continually pre-trained model instead of the fine-tuned model. We leverage Spearman’s rank correlation coefficient to measure the quality of all models. Among the baseline systems, the SimCSE-BERTLarge obtain higher results than other baseline approaches, but our presented method steadily outperforms all the baseline models.